OpenStudy (anonymous):

Find the value of y when x equals -4 in this equation. 5x + (-3y) = 13

OpenStudy (anonymous):

5x + (-3y) = 13 5(-4) -3y = 13 -20 -3y = 13 -3y = 13+20 -3y = 33 y = 33/-3 y = -11

OpenStudy (anonymous):

-11